A Bibliography of Publications in International Journal of High Speed Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.13

Title word cross-reference

0and1 [CC92b]. 3 [KY95]. 3/4 [HKC^{+}99]. A^{T} x [Ven94]. Ax [Ven94]. K [KR93, Ng00, AAD97, ASDOK04, DS96, MSAOK00]. M^{2}P = O(2^{nand2}) [CC92a]. n [AAD97, ASDOK04, DS96, MSAOK00]. s [PS93]. t [PS93].

-ary [AAD97, ASDOK04, DS96, MSAOK00]. -Cube [AAD97, ASDOK04, MSAOK00, DS96]. -Means-Type [Ng00]. -out-of- [PS93]. -Rotated [KR93].

100 [BCM^{+}93].

2 [DL90, EJL90, NM97]. 21st [Joh97].

3 [Pet94].

Abstract [NM92]. Access [Rin97]. Accessible [PMC90]. Achieving [Lee92a]. Activities [PA97]. Adaptive [CYM95, RG94, Shi00]. Address [Sat99]. Adjoints [BBW04]. Adviser [LZ96]. Affinity [WC95]. Agreement [CC95]. AI [Wal93]. Algebra [ADDM95]. Algorithm [ASDOK04, CC92a, CC92b, CC95, HLP95, Jen99, LC94, MM97, Nan94, NM97, PKL90, PY96, Shi00, SBK^{+}90, SJ95, TB95, TVM96, WC95, BJWS96, DS96]. Algorithms [ALA^{+}97, Cha94, CL92, DRS94, Fer95, GJM^{+}91, GM97c, GM97b, KHG96, Lee95, LL00, Lin99c, Ng00, PS90, SD94, THL^{+}96].

Allocation [AS95, KKY^{+}92, TVM96, TSKV00, VT96]. Almost [SJ95]. Alternate [PP99]. Ames [Bai93]. Analysis [BD92, CC92c, HSP^{+}99, LJ00, LS00, Lin99b, MSAOK00, PY96]. AND-parallel [NM92]. Annealing [MRP94, RP90]. Annealing-Based [MRP94, RP90]. Antennas [HSP^{+}99]. APE [BCM^{+}93]. APE-100 [BCM^{+}93]. Application [SS95]. Applications [HPK95, OSZ94, SBH96, Wil93]. Applied [ALMS92]. Approach [BASS95, PR95]. Approaches [RG93]. Approximate [CC95, KY95]. Arbitrary [CC92c]. Arbitrary-Delay [CC92c]. Architecture [BCM^{+}93, BD92, DO95, HL99, Joh97, RP90]. Architectures [AB93, Cia94, GP97, KRS96, PB89, WG94, YG96]. Arguments [Vij95]. Array [LS94]. Arrays [Lin00, SBK^{+}90, YC94]. ART1 [GP97]. ART2 [GP97]. Artificial [GP97]. ary [AAD97, ASDOK04, DS96, MSAOK00]. Assessment [Sch93]. Assignment [Ati99, LS94, MPT94]. assistant [CTS96]. ATM [BS99]. Author [Ano00]. Average [LS00, SKW92]. Average-Case [LS00].

Balanced [Lin99b, Lin99c, JC94a]. Balancing [BASS95]. Banyan [PYL^{+}94]. Barnes [BJWS96]. Barrier [LJ00]. Based [FK97, HL99, MRP94, MNP^{+}94, NM97, PC00, RP90, SCS90, TSKV00, FY92]. Benchmarking [Nag95]. Between [WG94]. Bidirectional [MM97]. Binary [MA90, PS95, SP92]. Block [BD89, Gir95]. Body [SRC^{+}95]. Bound [CC92c]. Boundary [BF92]. Bounded [YC94]. Bounds [SP92]. Broadcast [THL^{+}96, YC94]. Broadcasting [LHK97, LHK95]. Bus [HL99]. Bus-Based [HL99]. Butterfly [GH95].

Cache [MNP^{+}94, PC00, VZS99, WG94]. Cache-Based [MNP^{+}94]. Calculations [ADDM95]. Camera [Ano95a, Ano95b, Ano95c]. Camera-ready [Ano95a, Ano95b, Ano95c]. Carlo [SKG97, SBH90, SBH96, SWL90]. Case [Joh97, LS00, PB89, SKW92, WB95]. Casting [SS95]. Century [Joh97]. Channel [MRP94, THL^{+}96]. Characteristics [JP99]. Choice [CC92b]. Cholesky [MM97, RRS99, RG93]. Circuit [MSAOK00, RP90, RPM90]. Class [JC94b, NNN93, SC94]. Cluster [ES94]. Cluster-M [ES94]. Clusters [SWL90]. CM [NM97]. CM-2 [NM97]. Code [BF92]. coding [SCC96a, SCC96b]. Coefficient [GM97c, GM97b]. Coherence [PC00, WC94]. combinatorial [Li91]. Combinators [CC90]. Combining [Sat99]. Commercial [Joh97]. Communication [LL00, Lin99c, NLK99, OJH00, THL^{+}96, Swa04]. Comparative [BS99]. Comparison [SBH96]. Comparisons [Ma99, Ma04]. Compatible [HKC^{+}99]. Compilation [Nik93]. Compiler [EJL90]. Compilers [Cha94, CTS96]. complex [LPS96]. Complexity [Lin99b]. Compression [Jen99]. Computation [BBW04, DRS94, Fur93, SN00, Ven94]. Computational [NNN93, NM92]. Computations [Cha94, CC97, LS00]. Computer [ADDM95, BCM^{+}93, LW97, Ma04, Ng00, IFM90]. Computers [ALA^{+}97, Bai93, Erh90, GP94, HPK95, KY95, LT95, LHK95, SBH90, Wil93, WY90, YBN93, AS89]. Computing [BR93, BPS94, Joh97, LS94, OJH00, Rin97, TVM96, WB95, VT96]. Concepts [TM93]. Conflicts [MA90]. Conjugate [Gir95]. Connected [ALA^{+}97, LHK95]. Connection [JC94b, McB90, NM97, Sch93]. Connectivity [KHG96]. Consecutive [SS95]. Considering [TSKV00]. Consistency [WC99]. Constant [DRS94, GFNS99, KHG96]. Constrained [MNP^{+}94, VT96]. Constraints [PR95].

Contention [SC94]. Control [HL99]. Controls [SKY97]. Coordination [PA97]. Cost [Lin99b, LC97]. Cost/Performance [Lin99b]. CRAY [Ma99, CS95, DL90, EJL90]. Cray-2 [DL90]. CRAY-T3E [Ma99]. Cube [AAD97, ASDOK04, MSAOK00, DS96]. Cycle [SBS95]. Cycles [Lin99a, Lin00]. Cyclic [GM97a].

D [KY95]. Data [Cha94, GM97a, Jen99, JC94a, Kok93, MNP^{+}94, OSZ94, Rin97]. Data-balanced [JC94a]. Data-Parallel [Cha94]. Dataflow [AB93]. Decomposition [Cia94, BJWS96]. Decoupled [DO95]. deepening [Li91]. Definite [MM97]. Delay [CC92c]. Delayed [MA90]. Dense [AAD97]. Description [GP94]. Design [BL95, CTS96, LC95, PY96, RP90]. Development [LZ96]. Device [SKG97]. Diagonal [RR99]. Diagonal-Implicitly [RR99]. Difference [CC97, NNN93]. Differential [MR90]. Digital [DRS94]. Dimensional [DM90, LF96, SKG97]. Direct [GM97c, GM97b]. Discrete [SBH90, SBH96]. Distributed [ALMS92, Ati99, BR93, BD92, CC90, FGY95, Gir95, JC94a, LC94, MR92, Nan94, Ng00, PA97, RR99, TVM96, WC94, VT96]. Distributed-Memory [LC94, Nan94]. Distribution [GH95, LHK97, MNP^{+}94]. Divided [CC97]. Divided-Difference [CC97]. Division [HLQOSW00]. Domain [Cia94]. DSM [LJ00]. Dynamic [Ati99, LHK97]. Dynamics [NNN93, LPS96].

ECMWF [Hof93]. Effects [SC94]. Efficiency [LJ00]. Efficient [CS95, FK97, Fer95, HPK95, HS92, HLP95]. Eigenproblem [HLQOSW00]. Element [BL95]. Elimination [GM97a, MR92, NM97]. Elliptic [MR90]. Embedded [JC94b]. Embedding [Lin00].

Enhanced [PC00]. Environment [KS93]. Environments [BR93]. Equations [KRS96, MR90]. Equispaced [Vij95]. Evaluation [CC90, PS90, PS93, PB89, RP90, RG93, SKW92, Vij95]. Evolution [AB93]. Executable [LB93]. Executing [YG96]. Execution [LJ00, NM92, Zha92]. Experience [Bai93]. Experiences [Kar94]. Experiment [Zha92]. Exponentials [Vij95]. Extended [AS95]. Extensions [SBS95]. External [SP92]. Extremal [BPS94].

Factorisation [PP99]. Factorization [MM97, RRS99, RG93]. Factorized [KY95]. Fast [Jl97, NH00]. Fault [ASDOK04, JP99, PYL^{+}94, Shi00, Vij95]. Fault-Tolerant [ASDOK04, JP99, Shi00]. Faulty [Lin00]. FDTD [HSP^{+}99]. FE [KY95]. Feature [LT95]. Fewer [LHK95, KHG96]. Fibonacci [Pet94]. Fine [Dag93, Nic93]. Fine-Grain [Nic93]. Fine-Scale [Dag93]. Finite [NNN93]. Floating [Lee92b]. Flow [Kok93, Kuw92, SKY97]. Fluid [NNN93]. Forecasting [Hof93]. FORTRAN [DM90, CTS96, Nag95]. Four [DM90]. Four-Dimensional [DM90]. Free [CYM95]. Function [HLQOSW00, PKL90]. Functional [CC90, NM92]. Functions [Nag95].

GA [TSKV00]. Gauss [MR92]. Gaussian [NM97]. General [TB95]. General-Purpose [TB95]. Generalized [HLQOSW00, LHK95, Lin99b, Lin99c]. Generation [SBH90]. Generators [Pet94]. Genetic [Nan94, TVM96]. Geometry [DRS94]. Global [SD94]. Gradient [Gir95, JC94b]. Grain [Nic93]. Graph [KHG96, Lin99a]. Graphs [YG96]. Grouping [WC99].

Hamiltonian [Lin00]. Handling [SC94].

Hard [KS93]. Hardware [PC00]. HDG [LB93]. HDG-Machine [LB93]. Healing [BS99]. Hessenberg [BD89]. Heterogeneous [MPT94, PR95]. Heuristic [KKY^{+}92, MPT94]. Heuristics [AC92, SKW92]. Hierarchical [BJWS96, GFNS99, JP99, RG93, SCS90]. Hierarchical-Memory [RG93]. High [AS91, BL95, Lee92a]. Higher [MR90]. Histories [Kok93]. Homogeneous [ALMS92]. Hut [BJWS96]. Hybrid [CC90, WC94]. Hyper [THL^{+}96]. Hyper-Channel [THL^{+}96]. Hypercube [AS95, KKY^{+}92, Lee95, RP90, RG94, SS95, IFM90]. Hypercubes [CC97, GM97a, JP99, Lin99b, Lin99c, Mur90, Shi00].

i860 [Sin96]. i860^{T M} [Lee92a, Lee92b]. ICCG [HS92]. Id [Nik93]. Idle [SN00]. IEH [Lin99a]. II [GM97c, KY95]. Image [HPK95, PMC90]. Impact [BASS95]. Implementation [BD89, Cia94, FY92, GP97, PP99, RPM90, SCC96a, SCC96b, WY90, BJWS96, CTS96]. Implementations [DL90, HSP^{+}99]. Implementing [FGY95]. Implicit [NNN93]. Implicitly [RR99]. imprecise [Li91]. Improved [GM97a, Jen99]. Improving [GW95, LJ00]. Incompatible [Nic93]. Incompressible [LF96]. Incorporating [TVM96]. Index [Ano00]. Induction [PS90]. Informatics [WB95]. Information [Rin97]. Instruction [SC94, VZS99]. Instructions [Ano95a, Ano95b, Ano95c, CC92c]. Integer [Dag93]. Intel [SD94]. Interacting [SRC^{+}95]. Interactions [WG94]. Interconnected [BD92]. Interconnection [PYL^{+}94, PY96, DS96]. Interlocking [PP99]. Intermediate [GFNS99]. Intrinsic [Nag95]. Inverse [KY95]. IPSC [SD94]. IPSC/860 [SD94]. Irregular [Cha94]. Ising [DM90]. Isospeed [LS00]. Issue [VZS99]. Issues [Pol93]. Iterated [RR99].

Iteration [BD89]. Iterative

[DL90, JC94b, Mur90, SBK^{+}90, Li91].

iWarp [Gro93]. IWIM [PA97].

Kernel [LF96]. Knapsack

[CC92a, CC92b, Fer95, LC97]. Knowledge [TVM96]. Krylov [LW97]. Kutta [RR99].

Lagged [Pet94]. Language [NM92, Nik93]. Large [Gar93, Ma99, Ma04, Mur90]. LCAP [PB89]. Leapfrog [BBW04]. Learning [HL99]. Left [RG93]. Left-Looking [RG93]. Lempel [Jen99]. Length [BASS95]. LINDA [Kar94, FGY95]. Line [PMC90]. Linear [AAD97, ADDM95, GM97c, GM97b, KRS96, LC94, Lin00, Ma99, Ma04, Mur90, SCC96a, SCC96b]. Linked [PC00]. Linked-Based [PC00]. Lisp [FY92]. Load [Ati99, BASS95, LHK97, Sat99, TSKV00]. Local [MR90]. Locality [Rin97]. Logic [NM92, Zha92]. Looking [RG93]. Lopsided [SP92]. Lower [SP92].

M [ES94]. Machine

[JC94b, LB93, McB90, NM92, NM97,

SCC96a, SCC96b, Sch93, Swa04]. Machines [BF92, GW95, Nik93, RR99, RG93]. Manuscripts [Ano95a, Ano95b, Ano95c]. Many [SRC^{+}95]. Mapping [ALA^{+}97, AC92, CC97, GM97a, KRS96, YBN93]. Maps [LT95]. Marching [VvDP91]. market [LPS96]. MasPar [HPK95]. Massive [Fur93, Nic93]. Massively [BL95, Cia94, HPK95, KY95, LW97, Wal93, WY90]. Matrices [BPS94]. Matrix [Erh90, GM97c, GM97b, HLP95, HLQOSW00, Lee95, PKL90]. Matrix-Vector [HLP95]. Matter [TM93]. Means [Ng00]. Mechanism [PC00]. Medium [Dag93]. Memory [ADDM95, Ati99, BF92, BPS94, BD92, CC92a, EJL90, Fer95, Fur93, Gar93, Gir95, HS92, LC94, LC95, MR92, Nan94, Ng00, PMC90, RR99, RG93, WC94, WC95, WC99, YSA99].

Memory-Efficient [Fer95]. Memory/ Processor [CC92a]. Mesh [ALA^{+}97, GP97, Jl97, LHK95]. Mesh-Connected [LHK95]. Meshes [CC95, KHG96, LHK97, OSZ94]. Message [YG96]. Message-Passing [YG96]. Method [Cia94, LS94, PP99]. Methodology [PB89]. Methods [DL90, Gir95, LW97, MR90, RR99, SBH96]. Microprocessor [Lee92a, Lee92b, Sin96]. Migration [KKY^{+}92]. Migratory [PC00]. Millimeter [HSP^{+}99]. MIMD [Wil93, YBN93]. Minimal [WC95]. Model [DRS94, DM90, NM92, PA97, THL^{+}96, WY90, WC99, Zha92, IFM90, LPS96]. Modeling [DL94, PB89]. Modelled [PS93]. Models [Cha94]. Molecular [SWL90]. Monte [SKG97, SBH90, SBH96, SWL90]. Movement [OSZ94]. MP [CS95]. Multi [SKW92]. Multi-Processor [SKW92]. Multicast [FK97, PY96]. Multicomputer [CYM95]. Multifrontal [RG93]. Multigrid [LF96, SKG97, WY90]. Multiple [CC92b, LHK95, TSKV00, VZS99]. Multiple-Choice [CC92b]. Multiple-Issue [VZS99]. Multiplication [Erh90, HLP95, Lee95]. Multiprocessor [AS95, BR93, BPS94, DL94, Gar93, Gir95, GH95, HS92, KS93, PYL^{+}94, PS93, SCS90, Swa00, WC99, YSA99]. Multiprocessors [Ati99, Joh97, Lee95, LS00, LC94, MPT94, MR92, Nan94, RG94, WC95]. Multishift [BD89]. Multistage [PYL^{+}94, PY96, SCS90]. Multithreaded [DO95, WG94, WC99].

NASA [Bai93]. Naspack [Lee92a]. Navier [LF96]. NEC [Pet94]. Neocognitron [IFM90]. Network [CS95, JP99, OJH00, PYL^{+}94, SN00, IFM90, Kar94]. Networks [AAD97, ASDOK04, ALMS92, BS99, CYM95, GP97, MSAOK00, NLK99, PY96, SCS90, SKY97, DS96]. Neural [CS95, GP97, HL99, IFM90]. Non [GM97c, VZS99]. Non-Sequential [VZS99]. Non-Symmetric [GM97c]. Nondeterministic [Kok93]. Nonlinear [JC94b]. Nonsymmetric [Ma99, Ma04]. novel [DS96]. NOWP [FK97]. Number [Pet94]. Numerical [GJM^{+}91].

Objects [YSA99]. ODE [BF92]. onto [ALA^{+}97, GM97a, KRS96]. Optically [BD92]. Optimal [Lin99c, LC97, PS95]. Optimization [McB90, Sin96, Li91]. Optimizations [Rin97]. Optimizing [EJL90]. Order [MR90]. Oriented [MR92]. Overhead [OJH00, PC00, WC95]. Overview [Gro93].

P [AB93]. P-RISC [AB93]. p4 [SRC^{+}95]. Parabolic [VvDP91]. Paradigm [ES94]. Paralation [Sab93]. Parallel [AAD97, Bai93, BL95, BASS95, BF92, BBW04, CC92a, Cha94, Cia94, Dag93, DRS94, DL90, ES94, Far90, Fer95, Fur93, GJM^{+}91, GP94, GM97c, GM97b, HPK95, HLP95, Hof93, HL99, HSP^{+}99, HLQOSW00, Jl97, KY95, Lee95, Li91, LS94, LT95, LZ96, LS00, LC94, LF96, LW97, Ma99, Ma04, MRP94, MA90, Mur90, MM97, NH00, Nik93, NM97, PKL90, PA97, Pol93, PP99, PB89, RPM90, Rin97, SKG97, SWL90, SS95, SBK^{+}90, SJ95, TB95, THL^{+}96, VvDP91, Ven94, Wal93, WB95, WY90, WC99, AS89, BJWS96, IFM90, NM92, SCC96a, SCC96b, FY92]. Parallel-Multithreaded [WC99]. Parallel/Vector [Far90]. Parallelism [Dag93, Nic93]. Parallelization [NNN93]. Parallelizing [Nan94]. Parameters [BASS95]. Part [GM97c, GM97b]. Partial [MR90]. Partitioned [DL94]. Partitioner [RP90]. Passing [YG96]. Path [LL00, NLK99]. Pathlength [SP92]. PDE [JC94b]. 860 [SD94]. Performance [Lin99b]. Processor [CC92a]. Vector [Far90]. Performance [BL95, BASS95, BD92, GJM^{+}91, GW95, Lee92a, Lee92b, LZ96, McB90, MSAOK00, NH00, PY96, PS93, PB89, RP90, RG94, SBK^{+}90, VvDP91]. Periodic [KS93]. Pipelined [CC92b, CC92c, CC97, MSAOK00, SS95]. Planar [HSP^{+}99]. Point [Lee92b]. Polynomials [Vij95]. Positive [MM97]. Precedence [PR95, VT96]. Precedence-constrained [VT96]. Preconditioned [DL90, Gir95]. Preconditioners [Ma99, Ma04]. Preconditioning [KY95]. Prediction [GJM^{+}91]. Predictors [Sat99]. Prefetching [VZS99]. Prefix [SJ95]. Problem [AC92, CC92a, CC92b, Fer95, LC97, TVM96]. Problems [HL99, KHG96]. Processing [HPK95, Hof93, LC95, PMC90, SC94, WB95]. Processor [AS95, AS91, BL95, MPT94, SKW92]. Processors [CC92c, HKC^{+}99, KHG96, LHK95, PR95, VZS99]. Program [DM90, GFNS99, Zha92]. Programmable [TM93]. Programmer [BASS95]. Programming [BR93, Cha94, ES94, Gro93, Nik93, Pol93, Sab93]. Programs [BASS95, CC90, LZ96]. Propagation [GFNS99]. Properties [JP99]. Protocols [BS99, PC00]. Pseudo [BBW04]. Pseudo-Adjoints [BBW04]. Purpose [TB95].

QR [BD89]. Quadrant [PP99]. quadtree [SCC96a, SCC96b].

Radix [Jl97]. Random [Pet94, SBH90]. Randomized [AC92]. Rationale [Joh97]. Ray [SS95]. Read [MA90]. ready [Ano95a, Ano95b, Ano95c]. Real [Ati99, KS93, LL00, NLK99, SKY97]. Real-Time [Ati99, KS93, LL00, NLK99, SKY97]. Realization [IFM90, TM93]. reciprocals [Sin96]. Recognition [YSA99].

Recognizing [KR93]. Reconfigurable [Jl97, KHG96, OSZ94]. Recurrence [KRS96]. Recursive [PY96]. Redesigning [Far90]. Reduce [PC00]. Reduction [OJH00]. Redundancy [Vij95]. Refinement [MR90]. Register [EJL90]. Replicated [JC94a]. Replication [GM97a]. Representation [GFNS99]. Restriction [CYM95]. Restriction-Free [CYM95]. Right [RG93]. Right-Looking [RG93]. Ring [GP97, PYL^{+}94]. Ring-Banyan [PYL^{+}94]. Rings [Lin00]. RISC [AB93]. roots [Sin96]. Rotate [LHK95]. Rotated [KR93]. Router [MRP94]. Routing [ASDOK04, CYM95, RG94, Shi00, DS96]. Row [MR92, NM97]. Runge [RR99].

Sampling [SBH96]. Scalability [LS00, RRS99]. Scalar [SWL90]. Scale [Dag93, Gar93]. Scan [DRS94]. Scheduled [AS91, YG96]. Scheduling [CC92c, KS93, MNP^{+}94, PR95, SKW92, WC95]. Scheme [BBW04, Nan94, WC94, SCC96a, SCC96b]. Schemes [NNN93, VvDP91]. Scientific [Joh97, WB95, Wil93]. Search [JC94a, LC97, MA90, PR95]. Segments [KR93]. Selection [LL00, NLK99]. Self [BS99]. Semantics [Kok93]. Sequential [Jen99, VZS99]. Series [Pet94]. Shared [ADDM95, BF92, BPS94, BD92, Gar93, Gir95, HS92, WC94, WC95, WC99, YSA99]. Shared-Memory [BF92, BPS94, WC95, WC99]. Sharing [PC00]. Shrinking [SBS95]. Sign [HLQOSW00, PKL90]. SIMD [BL95]. Simple [Nan94, LPS96]. Simulated [MRP94]. Simulation [HKC^{+}99, Kuw92, Lin99a, RPM90, SKG97, Sat99, WG94]. Simulations [SBH90, SRC^{+}95, SWL90]. Singular [BPS94]. Skyline [Far90]. Software [LJ00, WC94]. Solid [YSA99]. Solution [AAD97, GM97c, GM97b, JC94b, KY95, Mur90, MM97]. Solver [Far90, LF96, SKG97]. Solving [AS89, LC94]. Some [DL90, GP94, Kar94]. Sort [Jl97, LHK95]. Sorting [CL92, Dag93, TB95, THL^{+}96]. Sparse [BPS94, Erh90, GM97c, GM97b, KY95, Ma04, MM97, RRS99, RG93, AS89]. Specific [TVM96]. Specification [LB93]. Spectral [HLQOSW00]. Speed [AS91]. square [Sin96]. Standard [VvDP91]. Star [PP99]. Static [AB93, LJ00, MPT94]. Statically [AS91]. stock [LPS96]. Stokes [LF96]. Strategy [Joh97, YBN93]. Streams [Kok93]. Structure [JC94a, SCS90]. Studies [SBK^{+}90]. Study [BS99, PB89, RG94, Sat99, WB95, WG94]. Subcube [KKY^{+}92]. Subroutines [Lee92a]. Subspace [LW97]. Suited [Cia94]. Sums [SS95]. Supercomputers [Far90, Joh97, Kuw92]. Supercomputing [ALMS92]. Supercube [Lin00]. Superscalar [AS91, LC95, SC94]. Switch [FK97]. Switch-Based [FK97]. Switching [MSAOK00]. SX [Pet94]. SX-3 [Pet94]. Symbolic [Fur93]. Symmetric [GM97c, GM97b, Joh97, MM97]. Synchronization [AS91, LJ00, SD94, WC95]. System [ALA^{+}97, BL95, BD92, FGY95, Gar93, Gro93, LC95, PMC90, PS93, YSA99]. Systems [AAD97, DL94, Fur93, GM97c, GM97b, KKY^{+}92, KY95, LS94, LC94, Ma99, Ma04, Mur90, MM97, PYL^{+}94, SRC^{+}95, SJ95, TVM96, WC99, AS89, VT96]. Systolic [KRS96, Vij95, YC94].

T3E [Ma99]. Tabu [PR95]. Task [BASS95, GH95, KKY^{+}92, PR95, TVM96, TSKV00, YG96, VT96]. Task-Length [BASS95]. Tasks [Ati99, KS93]. TC2000 [Gar93]. Technical [GP94]. Technique [LC97]. Techniques [EJL90, OSZ94]. Three [LF96, SKG97]. Three-Dimensional [LF96, SKG97]. Tiling [SN00]. Time [Ati99, BBW04, DRS94, KHG96, KS93, LL00, NLK99, SN00, SKY97, Vij95]. Time-Parallel [BBW04]. Toeplitz [SJ95]. Tolerant [ASDOK04, JP99, PYL^{+}94, Shi00, Vij95]. tool [CTS96]. Toolkit [HKC^{+}99]. Tools [BR93]. Topological [JP99, LT95]. Topology [PP99]. Traces [Kok93]. Tradeoff [CC92a]. Training [CS95]. Transformation [SN00]. Transforms [NH00]. Transputer [FY92, SBK^{+}90, Zha92]. Transputer-based [FY92]. Tree [GW95, SP92, BJWS96]. Trees [PS95]. triangular [AS89]. Tridiagonal [ALA^{+}97, LC94, SJ95]. Triplets [BPS94]. Tuplespace [FGY95]. Two [BS99, SKY97]. Type [Ng00]. Typed [CC92c]. Typesetting [Ano95a, Ano95b, Ano95c].

Upper [CC92c]. Usage [EJL90]. Using [GM97a, HLQOSW00, Jl97, Joh97, KHG96, LHK95, PY96, Rin97, SRC^{+}95, Vij95].

Value [BF92, Sat99]. Variables [SBH90]. Vector [DL90, Erh90, HLP95, HSP^{+}99, PS90, SBH90, Swa00, CTS96]. Vectorizable [SBH96]. Vectorization [CL92]. View [Hof93]. Virtual [ADDM95, Gir95, SCC96a, SCC96b]. Visualization [Kuw92]. Volume [Ano00]. vs [WB95].

Wave [HSP^{+}99]. Waveform [LW97]. Wavelet [NH00]. Weather [Hof93]. Web [OJH00]. Weighted [PS95]. Well [Cia94]. Well-Suited [Cia94]. Work [Fer95]. Workstations [SN00]. Wormhole [CYM95, FK97, NLK99, Shi00, SKY97]. Worst [SKW92]. Wraparound [CC95].

x86 [HKC^{+}99]. x86-Compatible [HKC^{+}99]. Xsim [HKC^{+}99].

Y-MP [CS95].

Ziv [Jen99].

References

Al-Ayyoub:1997:PSD

[AAD97] A-E Al-Ayyoub and K. Day.

Parallel solution of dense linear systems on the k-ary ncube networks. International Journal of High Speed Computing (IJHSC), 9(2):85–100, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Arvind:1993:EDA

[AB93] Arvind and S. Brobst. The evolution of dataflow architectures: From static dataflow to P-RISC. International Journal of High Speed Computing (IJHSC), 5(2):125–154, June 1993. CODEN IHSCEZ. ISSN 0129-0533.

Arunkumar:1992:RHM

[AC92] S. Arunkumar and T. Chockalingam. Randomized heuristics for the mapping problem. International Journal of High Speed Computing (IJHSC), 4 (4):289–300, December 1992. CODEN IHSCEZ. ISSN 0129-0533.

Amestoy:1995:LAC

[ADDM95] P. R. Amestoy, M. J. Daydé, Iain S. Duff, and P. Morère. Linear algebra calculations on a virtual shared memory computer. International Journal of High Speed Computing (IJHSC), 7(1):21–??, ????

1995. CODEN IHSCEZ. ISSN 0129-0533.

Amor:1997:MTS

[ALA^{+}97] M. Amor, J. Lopez, F. Arguello, et al. Mapping tridiagonal system algorithms onto mesh connected computers. International Journal of High Speed Computing (IJHSC), 9 (2):101–126, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Arbenz:1992:ADS

[ALMS92] P. Arbenz, H. P. Luthi, J. E. Mertz, and W. Scott. Applied distributed supercomputing in homogeneous networks. International Journal of High Speed Computing (IJHSC), 4(2):87– 108, June 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Anonymous:1995:ITCa

[Ano95a]  Anonymous.  Instructions

for typesetting camera-ready manuscripts. International Journal of High Speed Computing (IJHSC), 7(2):??, ???? 1995. CODEN IHSCEZ. ISSN 0129-0533.

Anonymous:1995:ITCb

[Ano95b]  Anonymous.  Instructions

for typesetting camera-ready manuscripts. International Journal of High Speed Computing (IJHSC), 7(3):??, ???? 1995. CODEN IHSCEZ. ISSN 0129-0533.

Anonymous:1995:ITCc

[Ano95c]  Anonymous.  Instructions

for typesetting camera-ready manuscripts. International Journal of High Speed Computing (IJHSC), 7(4):??, ???? 1995. CODEN IHSCEZ. ISSN 0129-0533.

Anonymous:2000:AIV

[Ano00]  Anonymous.  Author in-

dex: Volume 11 (2000). International Journal of High Speed Computing (IJHSC), 11 (4):253–254, December 2000. CODEN IHSCEZ. ISSN 0129-0533. URL http:

//ejournals.wspc.com.sg/
ijhsc/11/1104/S0129053300000199.
html.

Anderson:1989:SST

[AS89] E. C. Anderson and Y. Saad.

Solving sparse triangular systems on parallel computers. International Journal of High Speed Computing (IJHSC), 1 (??):73–96, 1989. CODEN IH-SCEZ. ISSN 0129-0533.

Arita:1991:HSS

[AS91] T. Arita and M. Sowa. High speed synchronization for a statically scheduled superscalar processor. International Journal of High Speed Computing (IJHSC), 3(1):77–??, March 1991. CODEN IHSCEZ. ISSN 0129-0533.

Ahuja:1995:PAE

[AS95] S. Ahuja and A. K. Sarje. Processor allocation in extended hypercube multiprocessor. International Journal of High Speed Computing (IJHSC), 7 (4):481–488, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Al-Sadi:2004:NFT

[ASDOK04] J. Al-Sadi, K. Day, and M. Ould-Khaoua. A new faulttolerant routing algorithm for k-ary n-cube networks. International Journal of High Speed Computing (IJHSC), 12(1):29– ??, June 2004. CODEN IH-SCEZ. ISSN 0129-0533.

Atif:1999:DLA

[Ati99] Y. Atif. Dynamic load assignment of real-time tasks in distributed memory multiprocessors. International Journal of High Speed Computing (IJHSC), 10(1):83–114, March 1999. CODEN IHSCEZ. ISSN 0129-0533.

Bailey:1993:EPC

[Bai93] D. H. Bailey. Experience with parallel computers at NASA Ames. International Journal of High Speed Computing (IJHSC), 5(1):51–62, March 1993. CODEN IHSCEZ. ISSN 0129-0533.

Ben-Asher:1995:LBP

[BASS95] Y. Ben-Asher, A. Schuster, and J. F. Sibeyn. Load balancing: A programmer’s approach or the impact of task-length parameters on the load balancing performance of parallel programs. International Journal of High Speed Computing (IJHSC), 7(2):303–??, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Bischof:2004:TPC

[BBW04] Christian H. Bischof, H. Martin Bücker, and Po-Ting Wu. Time-parallel computation of pseudo-adjoints for a leapfrog scheme. International Journal of High Speed Computing (IJHSC), 12(1):1–??, June 2004. CODEN IHSCEZ. ISSN 0129-0533.

Battista:1993:ACA

[BCM^{+}93] C. Battista, S. Cabasino, F. Marzano, et al. The APE-100 computer: (I) the architecture. International Journal of High Speed Computing (IJHSC), 5(4):637–656, December 1993. CODEN IH-SCEZ. ISSN 0129-0533.

Bai:1989:BIH

[BD89] Z. Bai and J. Demmel. On a block implementation of Hessenberg multishift QR iteration. International Journal of High Speed Computing (IJHSC), 1(1):97–112, 1989. CODEN IHSCEZ. ISSN 0129-0533. (also LAPACK Working Note #8).

Bogineni:1992:OID

[BD92] K. Bogineni and P. W. Dowd.

An optically interconnected distributed shared memory system: Architecture and performance analysis. International Journal of High Speed Computing (IJHSC), 4(3):179– 212, September 1992. CODEN IHSCEZ. ISSN 0129-0533.

Bennett:1992:PBV

[BF92] K. R. Bennett and G. Fairweather. A parallel boundary value ODE code for sharedmemory machines. International Journal of High Speed Computing (IJHSC), 4(2):71– 86, June 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Bhanot:1996:HDP

[BJWS96] B. Bhanot, J. Janak, R. Walkup, and V. Sonnad. Hierarchical decomposition: A parallel implementation of the Barnes– Hut tree algorithm. International Journal of High Speed Computing (IJHSC), 8(1):1– 12, ???? 1996. CODEN IH-SCEZ. ISSN 0129-0533.

Beal:1995:DPE

[BL95] D. Beal and C. Lambrinoudakis. Design of a processor element for a high performance massively parallel SIMD system. International Journal of High Speed Computing (IJHSC), 7(3):365–390, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Berry:1994:CES

[BPS94] M. W. Berry, B. N. Parlett, and A. H. Sameh. Computing extremal singular triplets of sparse matrices on a sharedmemory multiprocessor. International Journal of High Speed Computing (IJHSC), 6(2):239–276, 1994. CO-DEN IHSCEZ. ISSN 0129-

0533.  URL http://www.

worldscientific.com/doi/
abs/10.1142/S0129053394000123.

Bemmerl:1993:PTD

[BR93] T. Bemmerl and B. Ries.

Programming tools for distributed multiprocessor computing environments. International Journal of High Speed Computing (IJHSC), 5(4):595– 616, December 1993. CODEN IHSCEZ. ISSN 0129-0533.

Bajaj:1999:CST

[BS99] Vikas B. Bajaj and Anil K.

Sarje. A comparative study of two self healing protocols for ATM networks. International Journal of High Speed Computing (IJHSC), 10(3):235–256, September 1999. CODEN IH-SCEZ. ISSN 0129-0533.

Cheon:1990:HCD

[CC90] S. H. Cheon and J. W. Cho.

Hybrid combinators for distributed evaluation of functional programs. International Journal of High Speed Computing (IJHSC), 2(1):49– 68, March 1990. CODEN IH-SCEZ. ISSN 0129-0533.

Chang:1992:PAK

[CC92a] H. K.-C. Chang and J. J.- R. Chen. A. parallel algorithm for the knapsack problem with memory/processor tradeoff M^{2}P = O(2^{nand2}). International Journal of High Speed Computing (IJHSC), 4 (2):109–120, June 1992. CO-DEN IHSCEZ. ISSN 0129-

Chen:1992:PAM

[CC92b] G. H. Chen and M. S. Chern. A. pipelined algorithm for multiple-choice 0and1 knapsack problem. International Journal of High Speed Computing (IJHSC), 4(1):43–48, March 1992. CODEN IHSCEZ. ISSN 0129-0533.

Chou:1992:UBA

[CC92c] H.-C. Chou and C.-P. Chung. Upper bound analysis of scheduling arbitrary-delay instructions on typed pipelined processors. International Journal of High Speed Computing (IJHSC), 4(4):301–312, December 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Cheng:1995:AAA

[CC95] R. L. Cheng and C. P. Chung. An approximate agreement algorithm for wraparound meshes. International Journal of High Speed Computing (IJHSC), 7(3):407–420, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Chung:1997:MPD

[CC97] K-L Chung and Y-W Chen. Mapping pipelined divideddifference computations into hypercubes. International Journal of High Speed Computing (IJHSC), 9(3):181–190, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Chatterjee:1994:PMC

[Cha94] S. Chatterjee. Programming models, compilers, and algorithms for irregular dataparallel computations. International Journal of High Speed Computing (IJHSC), 6(2):183– 222, June 1994. CODEN IH-SCEZ. ISSN 0129-0533. URL

ftp://ftp.cs.unc.edu/pub/
users/sc/ijhsc-nested.PS.
Z.

Ciarlet:1994:IDD

[Cia94] P. Ciarlet. Implementation of a domain decomposition method well-suited for (massively) parallel architectures. International Journal of High Speed Computing (IJHSC), 6(1):157– ??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Chung:1992:VSA

[CL92] C.-P. Chung and W.-Y. Lin.

Vectorization of sorting algorithms. International Journal of High Speed Computing (IJHSC), 4(3):213–232, September 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Chung:1995:ENN

[CS95] S. L. Chung and R. Setiono.

Efficient neural network training on a Cray Y-MP. International Journal of High Speed Computing (IJHSC), 7(1):109– ??, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Chang:1996:DIF

[CTS96] C-Y Chang, J-Y Tzeng, and J-P Sheu. Design and implementation of a Fortran assistant tool for vector compilers. International Journal of High Speed Computing (IJHSC), 8 (1):13–46, 1996. CODEN IH-SCEZ. ISSN 0129-0533.

Chung:1995:RFA

[CYM95] J.-H. Chung, H. Yoon, and S. R. Maeng. Restrictionfree adaptive wormhole routing in multicomputer networks. International Journal of High Speed Computing (IJHSC), 7 (1):89–??, 1995. CODEN IH-SCEZ. ISSN 0129-0533.

Dagum:1993:PIS

[Dag93] L. Dagum. Parallel integer sorting with medium and finescale parallelism. International Journal of High Speed Computing (IJHSC), 5(4):503–522, December 1993. CODEN IH-SCEZ. ISSN 0129-0533.

Doi:1990:SPV

[DL90] S. Doi and A. Lichnewsky. Some parallel and vector implementations of preconditioned iterative methods on Cray-2. International Journal of High Speed Computing (IJHSC), 2(2):143–180, June 1990. CODEN IHSCEZ. ISSN 0129-0533.

Dowdy:1994:MPM

[DL94] L. W. Dowdy and M. R. Leuze. On modeling partitioned multiprocessor systems. International Journal of High Speed Computing (IJHSC), 6(1):31– ??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Drouffe:1990:FPF

[DM90] J.-M. Drouffe and K. J. M.

Moriarty. FORTRAN program for the four-dimensional Ising model. International Journal of High Speed Computing (IJHSC), 2(2):133–142, June 1990. CODEN IHSCEZ. ISSN 0129-0533.

Dorojevets:1995:MDA

[DO95] M. N. Dorojevets and V. G.

Oklobdzija. Multithreaded decoupled architecture. International Journal of High Speed Computing (IJHSC), 7(3):465– ??, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Djokic:1994:CTD

[DRS94] B. Djokic, J. Ruppert, and I. Stojmenovic. Constant time digital geometry algorithms on the scan model of parallel computation. International Journal of High Speed Computing (IJHSC), 6(4):501–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Deamine:1996:NRA

[DS96] E. D. Deamine and S. Srinivas. A. novel routing algorithm for k-ary n-cube interconnection networks. International Journal of High Speed Computing (IJHSC), 8(1):81–92, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Eisenbeis:1990:CTO

[EJL90] C. Eisenbeis, W. Jalby, and A. Lichnewsky. Compiler techniques for optimizing memory and register usage on the Cray 2. International Journal of High Speed Computing (IJHSC), 2(2):193–??, June 1990. CODEN IHSCEZ. ISSN 0129-0533.

Erhel:1990:SMM

[Erh90] J. Erhel. Sparse matrix multiplication on vector computers. International Journal of High Speed Computing (IJHSC), 2 (2):101–116, June 1990. CO-DEN IHSCEZ. ISSN 0129-0533.

Eshaghian:1994:CMP

[ES94] M. M. Eshaghian and M. E. Shaaban. Cluster-M parallel programming paradigm. International Journal of High Speed Computing (IJHSC), 6(2):287– ??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Farhat:1990:RSS

[Far90] C. Farhat. Redesigning the skyline solver for parallel/ vector supercomputers. International Journal of High Speed Computing (IJHSC), 2(3):223– 238, September 1990. CODEN IHSCEZ. ISSN 0129-0533.

Ferreira:1995:WME

[Fer95]  A. Ferreira.  Work and

memory-efficient parallel algorithms for the knapsack problem. International Journal of High Speed Computing (IJHSC), 7(4):595–606, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Feng:1995:ILT

[FGY95] M. D. Feng, Y. Q. Gao, and C. K. Yuen. Implementing Linda tuplespace on a distributed system. International Journal of High Speed Computing (IJHSC), 7(1):125–??, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Fan:1997:EMW

[FK97] K-P Fan and C-T King. Efficient multicast on wormhole switch-based NOWP. International Journal of High Speed Computing (IJHSC), 9(4):359– 380, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Furnari:1993:MSM

[Fur93] M. Furnari. Memory systems and massive parallel symbolic computation. International Journal of High Speed Computing (IJHSC), 5(3):307– 326, September 1993. CODEN IHSCEZ. ISSN 0129-0533.

Feng:1992:TBP

[FY92] M. D. Feng and C. K. Yuen. A transputer-based Parallel Lisp implementation. International Journal of High Speed Computing (IJHSC), 4(1):23–42, March 1992. CODEN IHSCEZ. ISSN 0129-0533.

Garber:1993:TSL

[Gar93] M. Garber. The TC2000 system — A large scale, shared memory, multiprocessor. International Journal of High Speed Computing (IJHSC), 5(3):475– 490, September 1993. CODEN IHSCEZ. ISSN 0129-0533.

Giordano:1999:CPH

[GFNS99] M. Giordano, M. M. Furnari, R. Napolitano, and A. Spagnolo. Constant propagation in a hierarchical intermediate program representation. International Journal of High Speed Computing (IJHSC), 10 (2):153–184, June 1999. CO-DEN IHSCEZ. ISSN 0129-0533.

Gottlieb:1995:TDB

[GH95] I. Gottlieb and A. Herold. Task distribution on a butterfly multiprocessor. International Journal of High Speed Computing (IJHSC), 7(1):1– ??, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Giraud:1995:BPC

[Gir95] L. Giraud. Block preconditioned conjugate gradient methods on a distributed virtual shared memory multiprocessor. International Journal of High Speed Computing (IJHSC), 7(2):161–190, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Gallivan:1991:PPP

[GJM^{+}91] K. Gallivan, W. Jalby, A. Malony, et al. Performance prediction for parallel numerical algorithms. International Journal of High Speed Computing (IJHSC), 3(1):31–62, March 1991. CODEN IHSCEZ. ISSN 0129-0533.

Gopalan:1997:IMC

[GM97a] K. Gopalan and C. Siva Ram Murthy. An improved mapping of cyclic elimination onto hypercubes using data replication. International Journal of High Speed Computing (IJHSC), 9(4):311–336, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Gopalan:1997:NPAb

[GM97b] K. Gopalan and C. Siva Ram Murthy. New parallel algorithms for direct solution of sparse linear systems: Part I — symmetric coefficient matrix. International Journal of High Speed Computing (IJHSC), 9 (4):259–290, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Gopalan:1997:NPAa

[GM97c] K. Gopalan and C. Siva Ram Murthy. New parallel algorithms for direct solution of sparse linear systems: Part II — non-symmetric coefficient matrix. International Journal of High Speed Computing (IJHSC), 9(4):291–310, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Gates:1994:TDS

[GP94] K. E. Gates and W. P. Petersen. A technical description of some parallel computers. International Journal of High Speed Computing (IJHSC), 6 (3):399–??, 1994. CODEN IH-SCEZ. ISSN 0129-0533.

Ghare:1997:IAA

[GP97] G. D. Ghare and L. M. Patnaik. Implementation of ART1 and ART2 artificial neural networks on ring and mesh architectures. International Journal of High Speed Computing (IJHSC), 9(1):41–56, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Gross:1993:OPI

[Gro93] T. Gross. An overview of programming the iWarp system. International Journal of High Speed Computing (IJHSC), 5 (3):379–402, September 1993. CODEN IHSCEZ. ISSN 0129-0533.

Gupta:1995:IPT

[GW95] A. K. Gupta and H. Wang. On improving the performance of tree machines. International Journal of High Speed Computing (IJHSC), 7(2):251–264, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Hsiao:1999:STX

[HKC^{+}99] Hung-Chang Hsiao, Chung-Ta King, Wei-Kuo Chen, Hsian-Hsiung Lin, and Chien-Chao Tseng. A simulation toolkit for x86-compatible processors 3/4 xsim. International Journal of High Speed Computing (IJHSC), 10(4):427–446, December 1999. CODEN IH-SCEZ. ISSN 0129-0533.

Hong:1999:PNL

[HL99] Tzung-Pei Hong and Jyh-Jong Lee. Parallel neural learning for control problems on a bus-based architecture. International Journal of High Speed Computing (IJHSC), 10 (3):257–274, September 1999. CODEN IHSCEZ. ISSN 0129-0533.

Hendrickson:1995:EPA

[HLP95] B. Hendrickson, R. Leland, and S. Plimpton. An efficient parallel algorithm for matrix-vector multiplication. International Journal of High Speed Computing (IJHSC), 7(1):73–??, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Huss-Lederman:2000:PSD

[HLQOSW00] Steven Huss-Lederman, Enrique S. Quintana-Ortí, Xiaobai Sun, and Yuan-Jye Y. Wu. Parallel spectral division using the matrix sign function for the generalized eigenproblem. International Journal of High Speed Computing (IJHSC), 11(1):1–14, March 2000. CODEN IHSCEZ. ISSN 0129-0533.

Hoffmann:1993:WFP

[Hof93] G.-R. Hoffmann. Weather forecasting and parallel processing: A view from ECMWF. International Journal of High Speed Computing (IJHSC), 5(1):63– 70, March 1993. CODEN IH-SCEZ. ISSN 0129-0533.

Hamdi:1995:EIP

[HPK95] M. Hamdi, Y. Pan, and W. T. Kwong. Efficient image processing applications on the MasPar massively parallel computers. International Journal of High Speed Computing (IJHSC), 7(4):489–514, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Hammond:1992:EIS

[HS92] S. W. Hammond and R. Schreiber. Efficient ICCG on a shared memory multiprocessor. International Journal of High Speed Computing (IJHSC), 4(1):1– 22, March 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Hoteit:1999:VPI

[HSP^{+}99] H. Hoteit, R. Sauleau, B. Philippe, Ph. Coquet, and J. P. Daniel. Vector and parallel implementations for the FDTD analysis of millimeter wave planar antennas. International Journal of High Speed Computing (IJHSC), 10(2):209–234, June 1999. CODEN IHSCEZ. ISSN 0129-0533.

Ito:1990:RNN

[IFM90] Takayuki Ito, Kunihiko Fukushima, and Sei Miyake. Realization of a neural network model Neocognitron on a hypercube parallel computer. International Journal of High Speed Computing (IJHSC), 2(1):1– 16, March 1990. CODEN IH-SCEZ. ISSN 0129-0533.

Johnson:1994:DRD

[JC94a] T. Johnson and A. Colbrook.

A distributed, replicated, databalanced search structure. International Journal of High Speed Computing (IJHSC), 6 (4):475–??, 1994. CODEN IH-SCEZ. ISSN 0129-0533.

Joubert:1994:EGI

[JC94b] W. D. Joubert and G. F.

Carey. Embedded gradient iterative solution of a class of nonlinear PDE’s on the connection machine. International Journal of High Speed Computing (IJHSC), 6(2):277–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Jena:1999:ILZ

[Jen99] S. K. Jena. An improved Lempel–Ziv algorithm for sequential data compression. International Journal of High Speed Computing (IJHSC), 10 (3):275–284, September 1999. CODEN IHSCEZ. ISSN 0129-0533.

Jang:1997:FPR

[Jl97] J-W Jang and K. G. lee. Fast parallel radix sort using a reconfigurable mesh. International Journal of High Speed Computing (IJHSC), 9(1):25– 40, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Johnston:1997:RSC

[Joh97] W. W. Johnston. Rationale and strategy for a 21st Century scientific computing architecture: The case for using commercial symmetric multiprocessors as supercomputers. International Journal of High Speed Computing (IJHSC), 9 (3):191–222, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Jayadevan:1999:FTC

[JP99] A. Jayadevan and L. M. Patnaik. Fault-tolerant characteristics and topological properties of a hierarchical network of hypercubes. International Journal of High Speed Computing (IJHSC), 10(1):1– 18, March 1999. CODEN IH-SCEZ. ISSN 0129-0533.

Karp:1994:SEN

[Kar94] A. H. Karp. Some experiences with Network LINDA. International Journal of High Speed Computing (IJHSC), 6(1):55– ??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Kao:1996:CTA

[KHG96] T-W Kao, S-J Horng, and Y-H Guo. Constant time algorithms for graph connectivity problems on reconfigurable meshes using fewer processors. International Journal of High Speed Computing (IJHSC), 8(4):371– 386, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Kang:1992:HSA

[KKY^{+}92] O. Kang, B. M. Kim, H. Yoon, et al. Heuristic subcube allocation with task migration in hypercube systems. International Journal of High Speed Computing (IJHSC), 4(2):121– 142, June 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Kok:1993:THS

[Kok93] J. N. Kok. Traces, histories and streams in the semantics of nondeterministic data flow. International Journal of High Speed Computing (IJHSC), 5 (2):225–242, June 1993. CO-DEN IHSCEZ. ISSN 0129-0533.

Katoen:1993:RRS

[KR93] J. P. Katoen and M. Rem. Recognizing K-rotated segments. International Journal of High Speed Computing (IJHSC), 5 (2):293–??, June 1993. CO-DEN IHSCEZ. ISSN 0129-0533.

Kazerouni:1996:MLR

[KRS96] L. Kazerouni, B. Rajan, and R. K. Shyamasundar. Mapping linear recurrence equations onto systolic architectures. International Journal of High Speed Computing (IJHSC), 8(3):229–270, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Khemka:1993:MSP

[KS93] A. Khemka and R. K. Shyamasundar. Multiprocessor scheduling of periodic tasks in a hard real-time environment. International Journal of High Speed Computing (IJHSC), 5 (4):617–636, December 1993. CODEN IHSCEZ. ISSN 0129-

Kuwahara:1992:FSS

[Kuw92] K. Kuwahara. Flow simulation on supercomputers and its visualization. International Journal of High Speed Computing (IJHSC), 4(1):49– ??, March 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Kolotilina:1995:FSA

[KY95] L. Yu. Kolotilina and A. Yu. Yeremin. Factorized sparse approximate inverse preconditioning II: Solution of 3D FE systems on massively parallel computers. International Journal of High Speed Computing (IJHSC), 7(2):191–216, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Lester:1993:ESH

[LB93] D. R. Lester and G. L. Burn. An executable specification of the HDG-machine. International Journal of High Speed Computing (IJHSC), 5(3):327– 378, September 1993. CODEN IHSCEZ. ISSN 0129-0533.

Lin:1994:PAS

[LC94] W.-Y. Lin and C.-L. Chen. A parallel algorithm for solving tridiagonal linear systems on distributed-memory multiprocessors. International Journal of High Speed Computing (IJHSC), 6(3):375–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Lu:1995:MSD

[LC95] N. P. Lu and C. P. Chung. Memory system design in superscalar processing. International Journal of High Speed Computing (IJHSC), 7(3):421– 444, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Lou:1997:COS

[LC97] D-C Lou and C-G Chang. A cost optimal search technique for the knapsack problem. International Journal of High Speed Computing (IJHSC), 9 (1):1–12, 1997. CODEN IH-SCEZ. ISSN 0129-0533.

Lee:1992:AHP

[Lee92a] K. Lee. Achieving high performance on the i860^{T M} microprocessor with Naspack subroutines. International Journal of High Speed Computing (IJHSC), 4(4):269–288, December 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Lee:1992:FPP

[Lee92b] K. Lee. On the floating point performance of the i860^{T M} microprocessor. International Journal of High Speed Computing (IJHSC), 4(4):251–268, December 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Lee:1995:PMM

[Lee95] P. Z. Lee. Parallel matrix multiplication algorithms on hypercube multiprocessors. International Journal of High Speed Computing (IJHSC), 7 (3):391–406, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Lou:1996:PTD

[LF96] J. Z. Lou and R. Ferraro. A parallel three-dimensional incompressible Navier–Stokes solver with a parallel multigrid kernel. International Journal of High Speed Computing (IJHSC), 8(4):319–346, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Lin:1995:GRS

[LHK95] C. F. Lin, S. J. Horng, and T. W. Kao. Generalized rotate sort on mesh-connected computers with multiple broadcasting using fewer processors. International Journal of High Speed Computing (IJHSC), 7 (4):515–530, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Lee:1997:DLD

[LHK97] W. Y. Lee, S. J. Hong, and J. Kim. Dynamic load distribution on meshes with broadcasting. International Journal of High Speed Computing (IJHSC), 9(4):337–358, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Li:1991:PII

[Li91] Tao Li. Parallel imprecise iterative deepening for combinatorial optimization. International Journal of High Speed Computing (IJHSC), 3(1):63– 76, March 1991. CODEN IH-SCEZ. ISSN 0129-0533.

Lin:1999:SCI

[Lin99a] Jen-Chih Lin. Simulation of cycles in the IEH graph. International Journal of High Speed Computing (IJHSC), 10 (3):327–342, September 1999. CODEN IHSCEZ. ISSN 0129-0533.

Lin:1999:BGHa

[Lin99b] Longsong Lin. Balanced generalized hypercubes: Complexity and cost/performance analysis. International Journal of High Speed Computing (IJHSC), 10 (4):379–398, December 1999. CODEN IHSCEZ. ISSN 0129-0533.

Lin:1999:BGHb

[Lin99c] Longsong Lin. Balanced generalized hypercubes: Optimal communication algorithms. International Journal of High Speed Computing (IJHSC), 10 (4):399–426, December 1999. CODEN IHSCEZ. ISSN 0129-0533.

Lin:2000:EHC

[Lin00]  Jen-Chih Lin.  Embed-

ding Hamiltonian cycles, linear arrays and rings in a faulty supercube. International Journal of High Speed Computing (IJHSC), 11(3): 189–201, September 2000. CODEN IHSCEZ. ISSN 0129-0533. URL http:

//ejournals.wspc.com.sg/
ijhsc/11/1103/S0129053300000151.
html.

Lee:2000:IEE

[LJ00] Jae Bum Lee and Chu Shik Jhon. Improving the execution efficiency of barrier synchronization in software DSM through static analysis. International Journal of High Speed Computing (IJHSC), 11 (3):167–188, September 2000. CODEN IHSCEZ. ISSN 0129-0533. URL http:

                                        //ejournals.wspc.com.sg/
                                        ijhsc/11/1103/S0129053300000138.
                                        html.

Lee:2000:PSA

[LL00] Yunkyung Lee and Sunggu Lee. Path selection algorithms for real-time communication. International Journal of High Speed Computing (IJHSC), 11 (4):215–222, December 2000. CODEN IHSCEZ. ISSN 0129-0533.

Levy:1996:CDS

[LPS96] M. Levy, N. Pesky, and S. Solomon. The complex dynamics of a simple stock market model. International Journal of High Speed Computing (IJHSC), 8(1):93–??, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Li:1994:AAM

[LS94] J. Z. Li and J. Srivastava. An array assignment method for parallel computing systems. International Journal of High Speed Computing (IJHSC), 6 (1):81–??, 1994. CODEN IH-SCEZ. ISSN 0129-0533.

Li:2000:ACA

[LS00] Keqin Li and Xian-He Sun. Average-case analysis of isospeed scalability of parallel computations on multiprocessors. International Journal of High Speed Computing (IJHSC), 11(1):15– 36, March 2000. CODEN IH-SCEZ. ISSN 0129-0533.

Li:1995:TFM

[LT95] Tao Li and Lixin Tao. Topological feature maps on parallel computers. International Journal of High Speed Computing (IJHSC), 7(4):531–546, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Luk:1997:WKS

[LW97] W-S Luk and O. Wing. Waveform Krylov subspace methods on a massively parallel computer. International Journal of High Speed Computing (IJHSC), 9(1):73–??, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Li:1996:PAD

[LZ96] K-C Li and K. Zhang. A performance adviser for the development of parallel programs. International Journal of High Speed Computing (IJHSC), 8 (3):205–228, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Meijer:1990:PBS

[MA90] H. Meijer and S. Akl. Parallel binary search with delayed read conflicts. International Journal of High Speed Computing (IJHSC), 2(1):17– 22, March 1990. CODEN IH-SCEZ. ISSN 0129-0533.

Ma:1999:CPP

[Ma99] Sangback Ma. Comparisons of the parallel preconditioners on the CRAY-T3E for large nonsymmetric linear systems. International Journal of High Speed Computing (IJHSC), 10 (3):285–300, September 1999. CODEN IHSCEZ. ISSN 0129-0533.

Ma:2004:CPP

[Ma04] Sangback Ma. Comparisons of the parallel preconditioners for large nonsymmetric sparse linear systems on a parallel computer. International Journal of High Speed Computing (IJHSC), 12(1):55–??, June 2004. CODEN IHSCEZ. ISSN 0129-0533.

McBryan:1990:OCM

[McB90] O. A. McBryan. Optimization of connection machine performance. International Journal of High Speed Computing (IJHSC), 2(1):23–48, March 1990. CODEN IHSCEZ. ISSN 0129-0533.

Murthy:1997:NBC

[MM97] K. N. Balasubramanya Murthy and C. Siva Ram Murthy. A new bidirectional Cholesky factorization algorithm for parallel solution of sparse symmetric positive definite systems. International Journal of High Speed Computing (IJHSC), 9 (1):57–72, 1997. CODEN IH-SCEZ. ISSN 0129-0533.

Min:1994:CBD

[MNP^{+}94] S. L. Min, J. H. Nam, M. S.

Park, et al. Cache-based data distribution constrained scheduling. International Journal of High Speed Computing (IJHSC), 6(1):139–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Menasce:1994:SHP

[MPT94] D. A. Menasce, S. C. Da Silva Porto, and S. K. Tripathi. Static heuristic processor assignment in heterogeneous multiprocessors. International Journal of High Speed Computing (IJHSC), 6(1):115–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

McCormick:1990:LRH

[MR90] S. F. McCormick and U. Rüde.

On local refinement higher order methods for elliptic partial differential equations. International Journal of High Speed Computing (IJHSC), 2(4):311– 334, December 1990. CO-DEN IHSCEZ. ISSN 0129-0533. Also available as TU-Bericht I-9034.

Mu:1992:ROG

[MR92] M. Mu and J. R. Rice. Row oriented Gauss elimination on distributed memory multiprocessors. International Journal of High Speed Computing (IJHSC), 4(2):143–168, June 1992. CODEN IHSCEZ. ISSN 0129-0533.

Mall:1994:PSA

[MRP94] R. Mall, S. Raman, and L. M. Patnaik. A parallel simulated annealing-based channel router. International Journal of High Speed Computing (IJHSC), 6(1):101–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Min:2000:PAA

[MSAOK00] Geyong Min, Hamid Sarbazi-Azad, and Mohamed Ould-Khaoua. Performance analysis of k-ary n-cube networks with pipelined circuit switching. International Journal of High Speed Computing (IJHSC), 11 (2):111–127, June 2000. CO-DEN IHSCEZ. ISSN 0129-0533.

Murthy:1990:PIS

[Mur90] C. Siva Ram Murthy. Parallel iterative solution of large linear systems on hypercubes. International Journal of High Speed Computing (IJHSC), 2(3):257– 264, September 1990. CODEN IHSCEZ. ISSN 0129-0533.

Nagai:1995:BFI

[Nag95] T. Nagai. Benchmarking Fortran intrinsic functions. International Journal of High Speed Computing (IJHSC), 7(2):217– 230, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Nang:1994:SPS

[Nan94] J. Nang. A simple parallelizing scheme of genetic algorithm on distributed-memory multiprocessors. International Journal of High Speed Computing (IJHSC), 6(3):451–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Ng:2000:MTA

[Ng00] Michael K. Ng. K-means-type algorithms on distributed memory computer. International Journal of High Speed Computing (IJHSC), 11(2):75– 92, June 2000. CODEN IH-SCEZ. ISSN 0129-0533.

Nielsen:2000:PPF

[NH00] Ole Møller Nielsen and Markus Hegland. Parallel performance of fast wavelet transforms. International Journal of High Speed Computing (IJHSC), 11 (1):55–74, March 2000. CO-DEN IHSCEZ. ISSN 0129-0533.

Nicolau:1993:MPF

[Nic93] A. Nicolau. Massive parallelism and fine-grain parallelism: Are they incompatible? International Journal of High Speed Computing (IJHSC), 5 (2):271–292, June 1993. CO-DEN IHSCEZ. ISSN 0129-0533.

Nikhil:1993:PPL

[Nik93] R. S. Nikhil. The parallel programming language Id and its compilation for parallel machines. International Journal of High Speed Computing (IJHSC), 5(2):171–223 (or 171–224??), June 1993. CO-DEN IHSCEZ. ISSN 0129-0533.

Nam:1999:PSR

[NLK99] Kyungwan Nam, Sunggu Lee, and Jong Kim. Path selection for real-time communication in wormhole networks. International Journal of High Speed Computing (IJHSC), 10 (4):343–360, December 1999. CODEN IHSCEZ. ISSN 0129-0533.

Nang:1992:PCM

[NM92] J. H. Nang and S. R. Maeng. An AND-parallel computational model and its abstract machine for the parallel execution of functional logic language. International Journal of High Speed Computing (IJHSC), 4(4):313–331 (or 313–332??), December 1992. CODEN IHSCEZ. ISSN 0129-0533.

Noh:1997:RBP

[NM97] S. H. Noh and A-M Moon. A row based parallel Gaussian elimination algorithm for the Connection Machine CM-2. International Journal of High Speed Computing (IJHSC), 9 (1):13–24, 1997. CODEN IH-SCEZ. ISSN 0129-0533.

Naik:1993:PCI

[NNN93] N. H. Naik, V. K. Naik, and M. Nicoules. Parallelization of a class of implicit finite difference schemes in computational fluid dynamics. International Journal of High Speed Computing (IJHSC), 5(1):1–50, March 1993. CODEN IHSCEZ. ISSN 0129-0533.

Oh:2000:ORN

[OJH00] Juil Oh, Ju-Wook Jang, and Tack-Don Han. Overhead reduction in the network communication for Web computing. International Journal of High Speed Computing (IJHSC), 11 (2):93–110, June 2000. CO-DEN IHSCEZ. ISSN 0129-0533.

Olariu:1994:DMT

[OSZ94] S. Olariu, J. L. Schwing, and J. Y. Zhang. Data movement techniques on reconfigurable meshes, with applications. International Journal of High Speed Computing (IJHSC), 6 (2):311–??, 1994. CODEN IH-SCEZ. ISSN 0129-0533.

Papadopoulos:1997:CDP

[PA97] G. A. Papadopoulos and F. Arbab. Coordination of distributed and parallel activities in the IWIM model. International Journal of High Speed Computing (IJHSC), 9(2):127– 160, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Prost:1989:MMP

[PB89] J. Prost and M. Becker. Modeling methodology for performance evaluation of parallel architectures a case study, LCAP. International Journal of High Speed Computing (IJHSC), 1(4):563–601, December 1989. CODEN IH-SCEZ. ISSN 0129-0533.

Pean:2000:ELB

[PC00] Der-Lin Pean and Cheng Chen. Enhanced linkedbased cache coherence protocols with a hardware mechanism to reduce the migratory sharing overhead. International Journal of High Speed Computing (IJHSC), 11 (4):223–252, December 2000. CODEN IHSCEZ. ISSN 0129-0533. URL http:

//ejournals.wspc.com.sg/
ijhsc/11/1104/S0129053300000163.
html.

Petersen:1994:LFS

[Pet94] W. P. Petersen. Lagged Fibonacci series random number generators for the NEC SX-3. International Journal of High Speed Computing (IJHSC), 6 (3):387–??, 1994. CODEN IH-SCEZ. ISSN 0129-0533.

Pandey:1990:PAM

[PKL90] P. Pandey, C. Kenney, and A. J. Laub. A parallel algorithm for the matrix sign function. International Journal of High Speed Computing (IJHSC), 2(2):181–192, June 1990. CODEN IHSCEZ. ISSN 0129-0533.

Park:1990:LAM

[PMC90] J. W. Park, S. R. Maeng, and J. W. Cho. Line accessible memory system for image processing. International Journal of High Speed Computing (IJHSC), 2(4):375–??, December 1990. CODEN IHSCEZ. ISSN 0129-0533.

Polychronopoulos:1993:PPI

[Pol93] C. D. Polychronopoulos. Parallel programming issues. International Journal of High Speed Computing (IJHSC), 5(3):413– 474, September 1993. CODEN IHSCEZ. ISSN 0129-0533.

Prameela:1999:PIA

[PP99] B. Prameela and L. M. Patnaik. Parallel implementation of alternate quadrant interlocking factorisation method on star topology. International Journal of High Speed Computing (IJHSC), 10(4):361–378, December 1999. CODEN IH-SCEZ. ISSN 0129-0533.

Porto:1995:TSA

[PR95] S. C. S. Porto and C. C.

Ribeiro. A tabu search approach to task scheduling on heterogeneous processors under precedence constraints. International Journal of High Speed Computing (IJHSC), 7 (1):45–??, 1995. CODEN IH-SCEZ. ISSN 0129-0533.

Pearson:1990:VEI

[PS90] R. A. Pearson and P. E.

Stokes. Vector evaluation in induction algorithms. International Journal of High Speed Computing (IJHSC), 2(1):85– ??, March 1990. CODEN IH-SCEZ. ISSN 0129-0533.

Prasad:1993:PEM

[PS93] E. V. Prasad and A. K. Sarje. Performance evaluation of multiprocessor system modelled as t-out-of-s system. International Journal of High Speed Computing (IJHSC), 5(1):71– 88, March 1993. CODEN IH-SCEZ. ISSN 0129-0533.

Pradhan:1995:OWB

[PS95] J. Pradhan and C. V. Sastry. On optimal weighted binary trees. International Journal of High Speed Computing (IJHSC), 7(3):445–464, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Park:1996:DPA

[PY96] J. Park and H. Yoon. Design and performance analysis of multistage interconnection networks using a recursive multicast algorithm. International Journal of High Speed Computing (IJHSC), 8(4):347– 362, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Park:1994:RBN

[PYL^{+}94] J. Park, H. Yoon, H. Lee, et al. The ring-banyan network: A fault tolerant multistage interconnection network for multiprocessor systems. International Journal of High Speed Computing (IJHSC), 6(4):557– ??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Rothberg:1993:ELL

[RG93] E. Rothberg and A. Gupta. An evaluation of left-looking, right-looking and multifrontal approaches to sparse Cholesky factorization on hierarchicalmemory machines. International Journal of High Speed Computing (IJHSC), 5(4):537– 594, December 1993. CODEN IHSCEZ. ISSN 0129-0533.

Reeves:1994:ARH

[RG94] D. S. Reeves and E. F.

Gehringer. Adaptive routing for hypercube multiprocessors: A performance study. International Journal of High Speed Computing (IJHSC), 6(1):1– ??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Rinard:1997:LOP

[Rin97] M. C. Rinard. Locality optimizations for parallel computing using data access information. International Journal of High Speed Computing (IJHSC), 9(2):161–??, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Raman:1990:ABC

[RP90] S. Raman and L. M. Patnaik. An annealing-based circuit partitioner for hypercube architecture: Design and performance evaluation. International Journal of High Speed Computing (IJHSC), 2(1):49– 68, March 1990. CODEN IH-SCEZ. ISSN 0129-0533.

Raman:1990:PIC

[RPM90] S. Raman, L. M. Patnaik, and R. Mall. Parallel implementation of circuit simulation.

International Journal of High Speed Computing (IJHSC), 2 (4):351–374, December 1990. CODEN IHSCEZ. ISSN 0129-0533.

Rauber:1999:DII

[RR99] T. Rauber and G. Rünger. Diagonal-implicitly iterated Runge–Kutta methods on distributed memory machines. International Journal of High Speed Computing (IJHSC), 10 (2):185–208, June 1999. CO-DEN IHSCEZ. ISSN 0129-0533.

Rauber:1999:SSC

[RRS99] T. Rauber, G. Rünger, and C. Scholtes. Scalability of sparse Cholesky factorization. International Journal of High Speed Computing (IJHSC), 10 (1):19–52, March 1999. CO-DEN IHSCEZ. ISSN 0129-0533.

Sabot:1993:PP

[Sab93] G. W. Sabot. Paralation programming. International Journal of High Speed Computing (IJHSC), 5(2):243–270, June 1993. CODEN IHSCEZ. ISSN 0129-0533.

Sato:1999:SSC

[Sat99] Toshinori Sato. A simulation study of combining load value and address predictors. International Journal of High Speed Computing (IJHSC), 10 (3):301–326, September 1999. CODEN IHSCEZ. ISSN 0129-0533.

Sarno:1990:GDR

[SBH90] R. Sarno, V. C. Bhavsar, and E. M. A. Hussein. Generation of discrete random variables on vector computers for Monte Carlo simulations. International Journal of High Speed Computing (IJHSC), 2(4):335– 350, December 1990. CODEN IHSCEZ. ISSN 0129-0533.

Sarno:1996:CVD

[SBH96] R. Sarno, V. C. Bhavsar, and E. M. A. Hussein. A comparison of vectorizable discrete sampling methods in Monte Carlo applications. International Journal of High Speed Computing (IJHSC), 8(3):295– ??, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Srinivas:1990:SPP

[SBK^{+}90] S. Srinivas, A. Basu, K. G. Kumar, et al. Studies on the performance of a parallel iterative algorithm on transputer arrays. International Journal of High Speed Computing (IJHSC), 2(3):265–??, September 1990. CODEN IHSCEZ. ISSN 0129-0533.

Sethi:1995:ECS

[SBS95] A. Sethi, S. Biswas, and A. Sanyal. Extensions to cycle shrinking. International Journal of High Speed Computing (IJHSC), 7(2):265–284, 1995. CODEN IHSCEZ. ISSN 0129-

Shiau:1994:EHI

[SC94] Y.-H. Shiau and C.-P. Chung. Effects and handling of instruction class contention in superscalar processing. International Journal of High Speed Computing (IJHSC), 6(3):357– ??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Shyu:1996:ILQ

[SCC96a] Shyong Jian Shyu, H. K.- C. Chang, and K.-C. Chou. Implementation of a linear quadtree coding scheme on the parallel virtual machine. International Journal of High Speed Computing (IJHSC), 8(1):65– 79, March 1996. CODEN IH-SCEZ. ISSN 0129-0533.

Sjyu:1996:ILQ

[SCC96b] J. Sjyu, H. K-C Chang, and K-C Chou. Implementation of a linear quadtree coding scheme on the parallel virtual machine. International Journal of High Speed Computing (IJHSC), 8 (1):65–80, 1996. CODEN IH-SCEZ. ISSN 0129-0533.

Schreiber:1993:ACM

[Sch93] R. Schreiber. An assessment of the Connection Machine. International Journal of High Speed Computing (IJHSC), 5 (4):523–536, December 1993. CODEN IHSCEZ. ISSN 0129-0533.

Sheu:1990:HMS

[SCS90] J. P. Sheu, W. T. Chen, and H. M. Su. A hierarchical multiprocessor structure based on multistage networks. International Journal of High Speed Computing (IJHSC), 2(2):117– 132, June 1990. CODEN IH-SCEZ. ISSN 0129-0533.

Seidel:1994:GSA

[SD94] S. R. Seidel and M. A. Davis.

Global synchronization algorithms for the Intel IPSC/860. International Journal of High Speed Computing (IJHSC), 6 (4):537–??, 1994. CODEN IH-SCEZ. ISSN 0129-0533.

Shih:2000:AFT

[Shi00] Jau-Der Shih. An adaptive fault-tolerant wormhole routing algorithm for hypercubes. International Journal of High Speed Computing (IJHSC), 11 (3):151–166, September 2000. CODEN IHSCEZ. ISSN 0129-0533. URL http:

//ejournals.wspc.com.sg/
ijhsc/11/1103/S012905330000014X.
html.

Sinclair:1996:ORS

[Sin96] R. Sinclair. Optimization of reciprocals and square roots on the i860 microprocessor. International Journal of High Speed Computing (IJHSC), 8(1):57– 64, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Sun:1995:PPA

[SJ95] X-H Sun and R. D. Joslin. A parallel prefix algorithm for almost Toeplitz tridiagonal systems. International Journal of High Speed Computing (IJHSC), 7(4):547–576, 1995.

CODEN IHSCEZ. ISSN 0129-0533.

Sandalci:1997:TDM

[SKG97] C. K. Sandalci, C. K. Koc, and S. M. Goodnick. Threedimensional Monte Carlo device simulation with parallel multigrid solver. International Journal of High Speed Computing (IJHSC), 9(3):223–236, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Shu:1992:WAC

[SKW92] R. B. Shu, M. S. Kankanhalli, and W. G. Wang. Worst and average case evaluation of heuristics for multi-processor scheduling. International Journal of High Speed Computing (IJHSC), 4(4):333–344, December 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Song:1997:TRT

[SKY97] H. Song, B. Kwon, and J-Y Kim H. Yoon. Two realtime flow controls in wormhole networks. International Journal of High Speed Computing (IJHSC), 9(3):237–258, 1997. CODEN IHSCEZ. ISSN 0129-0533.

Sathe:2000:CTI

[SN00] S. R. Sathe and P. M. Nawghare. Computation time and idle time of tiling transformation on a network of workstations. International Journal of High Speed Computing (IJHSC), 11(3): 129–149, September 2000.

CODEN IHSCEZ. ISSN 0129-0533. URL http:

//ejournals.wspc.com.sg/
ijhsc/11/1103/S0129053300000126.
html.

Sastry:1992:LBE

[SP92] C. V. Sastry and J. Pradhan. Lower bounds to the external pathlength of a lopsided binary tree. International Journal of High Speed Computing (IJHSC), 4(3):169–178, September 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Scalettar:1995:SIM

[SRC^{+}95] R. T. Scalettar, K. J. Runge, J. Correa, et al. Simulations of interacting many body systems using p4. International Journal of High Speed Computing (IJHSC), 7(3):327–350, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Song:1995:PPP

[SS95] J. Song and R. Shu. Parallel and pipelined parallel consecutive sums on a hypercube with application to ray casting. International Journal of High Speed Computing (IJHSC), 7 (1):145–??, 1995. CODEN IH-SCEZ. ISSN 0129-0533.

Swarztrauber:2000:VM

[Swa00] Paul N. Swarztrauber. The vector multiprocessor. International Journal of High Speed Computing (IJHSC), 11(1):37– 54, March 2000. CODEN IH-SCEZ. ISSN 0129-0533.

Swarztrauber:2004:CM

[Swa04] Paul N. Swarztrauber. The Communication Machine. International Journal of High Speed Computing (IJHSC), 12 (1):69–??, June 2004. CODEN IHSCEZ. ISSN 0129-0533.

Schutz:1990:MCS

[SWL90] M. Schutz, S. Wulfert, and S. Leutwyler. Monte Carlo simulations of molecular clusters: From scalar to parallel. International Journal of High Speed Computing (IJHSC), 2(4):289– 310, December 1990. CODEN IHSCEZ. ISSN 0129-0533.

Tridgell:1995:GPP

[TB95] A. Tridgell and R. P. Brent. A general-purpose parallel sorting algorithm. International Journal of High Speed Computing (IJHSC), 7(2):285–302, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Tsai:1996:PSA

[THL^{+}96] H-R Tsai, S-J Horng, S-S Lee, et al. Parallel sorting algorithms on a hyperchannel broadcast communication model. International Journal of High Speed Computing (IJHSC), 8(4):307–318, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Toffoli:1993:PMC

[TM93] T. Toffoli and N. Margolus. Programmable matter: Concepts and realization. International Journal of High Speed Computing (IJHSC), 5(2):155– 170, June 1993. CODEN IH-SCEZ. ISSN 0129-0533.

Tripathi:2000:GBM

[TSKV00] Anil Kumar Tripathi, Biplab Kumer Sarker, Naveen Kumar, and Deo Prakash Vidyarthi. A GA based multiple task allocation considering load. International Journal of High Speed Computing (IJHSC), 11(4):203–214, December 2000. CODEN IH-SCEZ. ISSN 0129-0533.

Tripathi:1996:GTA

[TVM96] A. K. Tripathi, D. P. Vidyarthi, and A. N. Mantri. A genetic task allocation algorithm for distributed computing systems incorporating problem specific knowledge. International Journal of High Speed Computing (IJHSC), 8(4):363– 370, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Venkatakrishnan:1994:PCA

[Ven94] V. Venkatakrishnan. Parallel computation of Ax and A^{T} x. International Journal of High Speed Computing (IJHSC), 6 (2):325–??, 1994. CODEN IH-SCEZ. ISSN 0129-0533.

Vijay:1995:FTS

[Vij95] M. Vijay. Fault tolerant systolic evaluation of polynomials and exponentials of polynomials for equispaced arguments using time redundancy. International Journal of High Speed Computing (IJHSC), 7(3):351– 364, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Vidyarthi:1996:PCT

[VT96] D. P. Vidyarthi and A. K. Tripathi. Precedence-constrained task allocation in distributed computing systems. International Journal of High Speed Computing (IJHSC), 8(1):47– 56, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Vandewalle:1991:PPS

[VvDP91] S. Vandewalle, R. van Driessche, and R. Piessens. The parallel performance of standard parabolic marching schemes. International Journal of High Speed Computing (IJHSC), 3 (1):1–30, March 1991. CODEN IHSCEZ. ISSN 0129-0533.

Veidenbaum:1999:NSI

[VZS99] A. V. Veidenbaum, Q. Zhao, and A. Shameer. Nonsequential instruction cache prefetching for multiple-issue processors. International Journal of High Speed Computing (IJHSC), 10(1):115–140, March 1999. CODEN IHSCEZ. ISSN 0129-0533.

Waltz:1993:MPA

[Wal93] D. L. Waltz. Massively parallel AI. International Journal of High Speed Computing (IJHSC), 5(3):491–??, September 1993. CODEN IHSCEZ. ISSN 0129-0533.

Waser:1995:CSP

[WB95] S. Waser and H. Burkhart.

A case study of parallel processing: Informatics vs scientific computing. International Journal of High Speed Computing (IJHSC), 7(4):577–594, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Wang:1994:HCS

[WC94] H.-H. Wang and R.-C. Chang.

A hybrid coherence scheme for software distributed shared memory. International Journal of High Speed Computing (IJHSC), 6(4):519–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Wang:1995:MSO

[WC95] Y. M. Wang and R. C. Chang.

A minimal synchronization overhead affinity scheduling algorithm for shared-memory multiprocessors. International Journal of High Speed Computing (IJHSC), 7(2):231–250, 1995. CODEN IHSCEZ. ISSN 0129-0533.

Wu:1999:GMC

[WC99] C.-C. Wu and C. Chen.

Grouping memory consistency model for parallelmultithreaded shared-memory multiprocessor systems. International Journal of High Speed Computing (IJHSC), 10(1):53– 82, March 1999. CODEN IH-SCEZ. ISSN 0129-0533.

Wong:1994:SSI

[WG94] W. F. Wong and E. Goto. A simulation study on the interactions between multithreaded architectures and the cache. International Journal of High Speed Computing (IJHSC), 6 (2):343–??, 1994. CODEN IH-SCEZ. ISSN 0129-0533.

Wilcke:1993:MCS

[Wil93] W. Wilcke. MIMD computers for scientific applications. International Journal of High Speed Computing (IJHSC), 5 (3):403–412, September 1993. CODEN IHSCEZ. ISSN 0129-0533.

Womble:1990:MIM

[WY90] D. E. Womble and B. C. Young. A model and implementation of multigrid for massively parallel computers. International Journal of High Speed Computing (IJHSC), 2 (3):239–256, September 1990. CODEN IHSCEZ. ISSN 0129-0533.

Yang:1993:MSM

[YBN93] J. Y. Yang, L. Bic, and A. Nicolau. A mapping strategy for MIMD computers. International Journal of High Speed Computing (IJHSC), 5 (1):89–??, March 1993. CO-DEN IHSCEZ. ISSN 0129-0533.

Yaacoby:1994:BBS

[YC94] Y. Yaacoby and P. Cappello. Bounded broadcast in systolic arrays. International Journal of High Speed Computing (IJHSC), 6(2):223–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Yang:1996:EST

[YG96] T. Yang and A. Gerasoulis. Executing scheduled task graphs on message-passing architectures. International Journal of High Speed Computing (IJHSC), 8(3):271–294, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Yaqub:1999:SMM

[YSA99] M. Yaqub, Q. Shaikh, and S. S. Ahmad. A shared memory multiprocessor system for the recognition of solid objects. International Journal of High Speed Computing (IJHSC), 10 (2):141–152, June 1999. CO-DEN IHSCEZ. ISSN 0129-0533.

Zhang:1992:ELP

[Zha92] K. Zhang. An experiment with a logic program execution model on the transputer. International Journal of High Speed Computing (IJHSC), 4(3):233– ??, September 1992. CODEN IHSCEZ. ISSN 0129-0533.